The fastest pitch recorded in baseball was 105.1 MPH (46.98 m/s) by Aroldis Chapman. Assuming the baseball had a mass of 0.145 kg, what was the impulse he imparted on the baseball?

If one was to catch a baseball moving at such a speed with their barehand (assume the catch takes place over a time interval of 0.020 seconds), what force would the baseball impart onto the hand? If hand bones can break when under ~110 Newtons of force, can one safely catch such a ball?

please answer both of these questions and their parts as they are related
Answer #1

we know that,

Change in momentum(P) = impulse imparted(I)

m(v-u) = F*t

m= 0.145 kg

When the ball was first thrown,

u = 0, v = 46.98 m/s,

I = m(v-u)

I = 0.145*(46.98 - 0)

I = 6.8121 N-s

Now again,

u = 46.98 m/s, v = 0 m/s,

I = m(v-u)

F*0.02 = 0.145*(0 - 46.98)

F = -340.605 N

-ve only tells the direction

If the ball is caught in the above specified time period catching it would be fatal,

One other way to safely catch the ball is to take more time to catch the ball, i.e, pull your hands backwards as you catch the ball